### v2

# Interactive Query Answering on Knowledge Graphs with Soft Entity Constraints

Abstract
Methods for query answering over incomplete knowledge graphs retrieve entities that are likely to be answers, which is particularly useful when such answers cannot be reached by direct graph traversal due to missing edges. However, existing approaches have focused on queries formalized using first-order-logic. In practice, many real-world queries involve constraints that are inherently vague or context-dependent, such as preferences for attributes or related categories. Addressing this gap, we introduce the problem of query answering with soft constraints. We formalize the problem and introduce two efficient methods designed to adjust query answer scores by incorporating soft constraints without disrupting the original answers to a query. These methods are lightweight, requiring tuning only two parameters or a small neural network trained to capture soft constraints while maintaining the original ranking structure. To evaluate the task, we extend existing QA benchmarks by generating datasets with soft constraints. Our experiments demonstrate that our methods can capture soft constraints while maintaining robust query answering performance and adding very little overhead.
1 Introduction
Knowledge graphs (KGs) are graph-structured databases that store information about a domain using triples of the form (subject, predicate, object). Such a compact representation enables efficient algorithms for answering queries about entities over the domain encoded by the KG (fensel_knowledge_2020; hogan_knowledge_2021). In principle, answers to these queries can be obtained by traversing the graph and collecting the entities that satisfy the given conditions. However, this strategy misses answers due to connections that are not explicit in a KG that may be incomplete, but that are likely to satisfy a query, for example because of their similarity to known answers. The problem of finding likely answers to simple queries over KGs has been studied in areas such as rule mining (galarraga_amie_2013; meilicke_fine-grained_2018; meilicke_anytime_2019; qu_rnnlogic_2020) and link prediction (nickel_review_2015; ji_survey_2020), and complex query answering (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural). Given a query, these methods score all the entities in the KG, indicating the likelihood of meeting all conditions specified in the query. Notably, prior works have focused on conditions that can be specified using different subsets of first order logic.
We identify a relevant and complementary problem: answering queries with constraints that cannot be specified formally as further conditions in a query. We refer to these as soft constraints, to contrast them with the hard constraints that can be specified using first order logic. Consider the following query: “What are award nominations received by movies starring Leonardo DiCaprio?”. A user might be additionally interested in nominations “related to the audio-visual design of the movie”, such as Best Costume Design, or Best Sound, and not in nominations about the artistic vision of a film (see 1). Such a constraint cannot be specified using the language of first order logic, simply because such an idea of relatedness is too vague to be formalized into precise logical statements. Our goal is thus to broaden the expressivity of queries on KGs which are specified using first order logic, by incorporating additional soft constraints to refine the set of answers.
Figure 1: A query over a KG can be written in a logical form that specifies hard constraints that an entity must meet. We investigate the problem of incorporating soft constraints that cannot be specified in logical form, such as “awards given to performers.” Such constraints correspond to subsets of entities, shown in shaded clusters, that share some similarities.
The challenge of incorporating soft constraints lies in adjusting answer scores with minimal deviation from the original distribution of scores, so that hard logical constraints imposed by the query are still met. Training a model for this task requires datasets that capture the relationships between query answers and soft constraints – data that is not available in existing query answering benchmarks over incomplete KGs. Additionally, the reranking process must be efficient and integrate seamlessly with state-of-the-art QA models.
To tackle these challenges, we formalize the problem and introduce efficient methods designed to capture soft constraints while preserving the global structure in an original list of scores for a given query. In summary, our work makes the following key contributions:
  1. 1.
    We introduce the problem of query answering on knowledge graphs with soft entity constraints, defining it as an interactive process where preferences are specified incrementally.
  2. 2.
    We introduce two computationally efficient methods for reranking query answers based on an initial list of scores from an existing query answering system. The methods are lightweight, requiring tuning two parameters or a small neural network optimized to capture soft constraints while maintaining the original ranking structure.
  3. 3.
    We extend existing benchmarks for query answering with automatically generated data of soft constraints. Through comprehensive evaluations, we show how our methods effectively capture these constraints without significantly compromising global ranking performance or runtime performance.
2 Related Work
Approximate Query Answering.
We build on methods for learning on KGs that retrieve query answers that are not reachable via graph traversal. Early work focused on simple link prediction, where an edge is predicted between two entities (nickel2011rescal; bordes2013transe; trouillon2016complex; lacroix2018complex_n3; sun2019rotate). More recently, the set of queries that learning-based methods can tackle has been broadened to complex queries (e.g., paths of length two or three with intermediate variables), using a formulation of queries based on a small (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural) or a larger  (cucumides_unravl_2024; yin_rethinking_2024) subset of first order logic. In all these cases, the output of a query answering model is a score for each entity in the graph that indicates how likely it is to meet constraints specified using first order logic. Our work complements these approaches by supporting soft constraints that cannot be expressed in the same way, due to the inherent limitations of the language of first order logic.
Active Learning on KGs. The idea of collecting entities that exemplify soft constraints is closely related to the problem of active learning (fu_survey_2013; ren_survey_2022). Active learning allows training models incrementally when data is of limited availability. In the context of KGs, it has been applied to sampling training data for link prediction models under a constrained computational budget (ostapuk_activelink_2019), error detection (dong_active_2023), and KG alignment (huang_deep_2023); all cases in which a prediction is well-defined. Applying active learning directly to an existing query answering model in order to capture soft constraints is not trivial, as a single query can have different types of soft constraints (as illustrated in Fig. 1), potentially leading to contradicting model updates. We avoid this by training a lightweight neural query reranker via supervised learning, using automatically generated data of soft preferences.
Information Retrieval (IR) and Recommender Systems. A common task in IR is to retrieve documents relevant to a given query (manning2009introduction), for example based on their similarity to a textual query using sequence embeddings (kenton2019bert). We adopt a similar approach in one of the baselines in our experiments. We also draw inspiration from Learning to Rank (LTR), where models are trained to order items according to their relative relevance. Pairwise methods such as RankNet and LambdaRank (burges_learning_2005; burges_learning_2006) and their boosted extension LambdaMART (burges2010ranknet) learn from preference pairs by directly optimizing ranking metrics such as NDCG, and have been shown to outperform regression- or classification-based ranking approaches. Modern implementations, such as the one provided in LightGBM (ke2017lightgbm), efficiently extend these methods to large datasets while supporting graded relevance labels and fast tree-based inference. These approaches generate a new ranking from scratch given a set of preferences; we, on the other hand, aim to preserve the original score relationships while biasing the ranking towards soft constraints. Other learning-to-rank methods take a list-wise approach that considers permutations of candidate lists (cao_learning_2007; xia_listwise_2008; zhu_listwise_2020), but these approaches do not scale well to the number of entities typically found in a knowledge graph. Lastly, the field of Recommender Systems has explored the use of interaction data (e.g., clicks, ratings, and browsing history) to generate recommendations (koren2021advances). Various collaborative filtering models exist in the literature (ning2011slim; steck2021negative; koren2009matrix; rendle2012bpr; he2017neural; christoffel2015blockbusters), some of which demonstrate the effectiveness of neural networks at capturing preference data.
3 Preliminaries
We denote a KG as a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations, and $\mathcal{E}$ is the set of edges of the form $(h,r,t)$ where $h,t\in\mathcal{V}$ are the head and tail entities, and $r\in\mathcal{R}$ is the relation between them. A triples $(h,r,t)$ in a KG implies the truth of the statement $r(h,t)$ in first-order logic, where $r\in\mathcal{R}$ is a binary predicate, and $h,t\in\mathcal{V}$ its arguments.
Logical queries over KGs. A query $q$ over $\mathcal{G}$ defines a set of constraints that an entity must satisfy to be considered an answer. For example, consider the query What are award nominations received by movies starring Leonardo DiCaprio? The target variable of this query is an award, and the query places two constraints on it: 1) the award nomination is given to movies, and 2) the movie features Leonardo DiCaprio. Formally, we can express the set $A_{q}$ of answers to the query using first order logic as follows:
$A_{q}=\{v_{2}\mid\exists v_{1},v_{2}\in\mathcal{V}:\texttt{performer}(\texttt{Leonardo DiCaprio},v_{1})\land\texttt{nominated}(v_{1},v_{2})\}.$ (1)
Answering this query requires assigning entities from $\mathcal{V}$ to variables $v_{1}$ and $v_{2}$ such that the conjunction specifying the constraints is true according to the information contained in the KG. More generally, and following the notation in bai2023qto, we consider first-order logic (FOL) queries that can include conjunctions, disjunctions, and negations, which we write in disjunctive normal form as follows:
$\mathcal{A}_{q}=\{v_{t}\mid\exists v_{1},\dots,v_{N}\in\mathcal{V}:\,(c^{1}_{1}\land\dots\land c^{1}_{m_{1}})\lor\dots\lor(c^{n}_{1}\land\dots\land c^{n}_{m_{n}})\},$ (2)
where $v_{1},\ldots,v_{N}$ are variables in the query, and $v_{t}$ is the target variable. The constraints $c_{j}^{i}$ of the query consist of binary relations between two variables, or between a variable and a known entity in $\mathcal{V}$, while optionally including negations:
$c^{i}_{j}=\begin{cases}r(e,v)\text{ or }r(v^{\prime},v)\\ \lnot r(e,v)\text{ or }\lnot r(v^{\prime},v)\end{cases}\hskip 30.00005pt\text{with }e\in\mathcal{V},v,v^{\prime}\in\{v_{t},v_{1},\ldots,v_{N}\},\text{ and }r\in\mathcal{R}.$
Approximate Query Answering. In order to provide answers to queries that are not limited to the information explicitly stated by the edges in the graph, recent works have introduced machine learning models for approximate query answering (QA) (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022; ren2024neural; cucumides_unravl_2024; yin_rethinking_2024). For a given query, these methods compute a score for all the entities in the KG according to how likely they are to be answers to a query. While the specific mechanisms for achieving this vary among methods, in general we denote a QA model as a function $a$ that maps a query $q$ to a vector $a(q)\in\mathbb{R}^{|\mathcal{V}|}$ of real-valued scores, with one score for each entity in the KG.
4 Query Answering with Soft Constraints
FOL queries over KGs allow imposing a broad range of constraints over entities, but this approach is still limited by the expressivity of first-order logic and the schema of the KG. In some applications one might require incorporating notions of similarity, which can be hard or impossible to express using FOL. We address this limitation by introducing the concept of soft constraints.
Definition 1 (Soft constraints (informal)).
A soft constraint is a loose requirement based on semantic similarity, which can guide answers to a query towards desirable types of answers without imposing strict logical criteria. For example, the soft constraint “awards involving performers” suggests prioritizing answers related to acting awards, without strictly excluding other awards.
We consider a scenario where soft constraints are specified by means of specific examples of entities that represent the desired (or undesired) type of answer. We formalize this through labeled preferences. A preference is a tuple $(e,l)$, where $e\in\mathcal{V}$ is an entity in the KG, and $l\in\{0,1\}$ is a label indicating whether entities similar to $e$ should be preferred ($l=1$) or de-emphasized ($l=0$).
Since specifying an exhaustive list of preferences can be impractical, we consider an interactive setting in which preferences are provided incrementally.
Definition 2 (Interactive query answering with soft constraints).
Let $q$ be a query with hard logical constraints evaluated over a knowledge graph, and let $a^{(0)}(q)$ denote the initial list of scores produced by a query answering model. In the interactive query answering with soft constraints setting, the scores are adjusted iteratively based on a growing set of user-provided preferences: after $t$ interactions, the preference set is
$P(t)=\{(e_{1},l_{1}),\ldots,(e_{t},l_{t})\}.$ (3)
At each iteration $t\geq 1$, we denote the adjusted list of scores as
$a^{(t)}(q)=f\big(a^{(0)}(q),P(t)\big),$ (4)
where $f$ is a function that updates the initial scores according to the accumulated preferences.
This formulation enables incorporating soft constraints dynamically, guiding the ranking of answers toward entities semantically aligned with the provided examples. Consider the query “What are award nominations received by movies starring Leonardo DiCaprio?” The corresponding logical form retrieves all entities $v_{2}$ such that there exists a movie $v_{1}$ with $\texttt{performer}(\texttt{Leonardo DiCaprio},v_{1})$ and $\texttt{nominated}(v_{1},v_{2})$. Answers to the hard constraints include awards such as Best Actor, Best Picture, and Best Sound. In the case where we are interested in nominations “related to the audio-visual design of the movie”, but not those “about the artistic vision”, we can express this through the preference set
$P(2)=\{(\texttt{Best Sound},1),(\texttt{Best Picture},0)\}.$ (5)
Here, the label $l=1$ marks Best Sound as a positive example (entities similar to it (e.g., Best Costume Design, Best Original Song) should receive higher scores) while $l=0$ marks Best Picture as a negative example, encouraging the model to de-emphasize similar entities (e.g., Best Director). After incorporating $P(2)$, the updated scores $a^{(2)}(q)$ should give priority to entities that reflect the soft constraint.
5 Incorporating Soft Constraints
We now turn our attention to the problem of adapting answer scores based on preference sets that represent a soft constraint. The goal is to develop a method that can dynamically update the ranking of answers in response to incremental feedback of preferences, while maintaining the structure of the initial scores provided by the base QA model. Following our running example, given a soft preference like “awards involving performers”, we would like awards such as “Best Actress” to rank higher than “Best Sound”, but both entities should still rank higher than incorrect entities, such as “Nobel Prize”.
To achieve this, we propose to modify the original score of each entity with a linear function with three inputs: (1) the original score computed by the base QA model, (2) its average similarity to positively labeled entities, and (3) its average similarity with negatively labeled entities. Formally, we start by defining a partition of a preference set into the following two sets:
$P^{+}=\{e\mid(e,l)\in P(t),\,l=1\},$ (6)
$P^{-}=\{e\mid(e,l)\in P(t),\,l=0\},$ (7)
where we have dropped the dependency on $t$ for brevity. Let $a[e]$ denote the score of entity $e$ extracted from the vector of scores $a(q)$ computed by the base QA model. Our proposed score update is given by the following expression:
$a^{\text{new}}[e]=f(a[e],\delta(P^{+},e),\delta(P^{-},e)),$ (8)
where $f$ is a linear function of its arguments and $\delta$ is a function that computes the mean cosine similarity of a preference set and an entity $e$:
$\delta(P^{\odot},e)=\frac{1}{|P^{\odot}|}\sum_{e_{i}\in P^{\odot}}\operatorname{sim}(e_{i},e).$ (9)
The majority of existing QA models learn entity embeddings for computing query scores (hamilton2018gqe; daza2020message; ren_query2box_2020; ren_beta_2020; arakelyan2021cqd; arakelyan2023cqdA; bai2023qto; zhu_neural-symbolic_2022, among others). We propose to reuse the embeddings of the base QA model (which are already required to compute the initial vector of scores $a(q)$) to compute the cosine similarity $\operatorname{sim}(e_{i},e)$ required in Equation˜9.
Our proposed score update has two benefits: First, the use of a linear function preserves monotonicity in the scores, such that if two entities have equal similarities with respect to a preference set, then differences in their scores will be based exclusively on the initial scores computed by the QA model. Second, it is efficient, since it relies on already trained embeddings from the base QA model, and its application has linear time complexity, as we show next.
Time complexity.
Let $n=|\mathcal{V}|$ be the number of entities, $d$ the embedding dimension, and $t=|P(t)|$ the number of preferences. For each entity $e\in\mathcal{V}$ we compute $\delta(P^{\odot},e)$ as the mean of $t$ cosine similarities. Each cosine reduces to a dot product between $d$-dimensional vectors, so it costs $\mathcal{O}(d)$; hence one entity requires $\mathcal{O}(kd)$ time to accumulate $p$ dot products and take their mean, and processing all $n$ entities costs $\mathcal{O}(nkd)$ time.
We now introduce two implementations of the update function $f$ in Equation˜8: a linear combination with fixed coefficients, and a neural reranker where the coefficients are computed conditionally.
5.1 Linear with fixed coefficients (Cosine).
Motivated by prior work on reranking methods for search over knowledge graphs (gerritse2020graph; daza2021blp), we first propose a score update that is computed as a convex combination of the original score, and a linear combination of the average cosine similarities with $P^{+}$ and $P^{-}$. We refer to this as the Cosine update:
$f_{\text{Cos}}(a[e],\delta(P^{+},e),\delta(P^{-},e))=\alpha\cdot a[e]+(1-\alpha)\left(\frac{1+\beta}{2}\delta(P^{+},e)-\frac{1-\beta}{2}\delta(P^{-},e)\right)$ (10)
The cosine update thus requires tuning two parameters, $\alpha\in(0,1)$ and $\beta\in(-1,1)$ that control the weight given to the original score, and the average cosine similarities, which can be done using a validation set.
5.2 Neural query reranker (NQR).
We can further adapt the weights given to each term in the score update function in Equation˜10 by computing them with a small multi-layer perceptron (MLP) with four inputs: the original score $a[e]$, the average similarities $\delta(P^{+},e)$ and $\delta(P^{-},e)$, and the embedding $\mathbf{e}\in\mathbb{R}^{d}$ of entity $e$ (which we reuse from the base QA model and do not fine-tune).
Intuitively, making the weights conditional allows to contextualize the score update to specific values of scores, similarities, and the entity as captured by its embedding. The MLP outputs two scalars $\alpha_{p},\alpha_{n}\in(0,1)$, which we then use in the update rule. We refer to this as the NQR update:
$f_{\text{NQR}}(a[e],\delta(P^{+},e),\delta(P^{-},e))=a[e]+\alpha_{p}\delta(P^{+},e)-\alpha_{n}\delta(P^{-},e).$ (11)
Time complexity.
If the hidden dimension of the MLP is $h$, the forward pass per entity multiplies a $(d+3)\times h$ matrix and an $h\times 2$ matrix (plus activations), which is $\mathcal{O}(dh)$ time; across all entities this adds $\mathcal{O}(ndh)$. Thus the NQR update over all entities runs in $\mathcal{O}(n(kd+dh))$ time, remaining linear in the number of entities in the KG.
Training NQR.
The main goal of NQR is to adjust entity scores such that the information conveyed by the preference data is reflected in the ranking, while preserving the overall quality of the original query answers. This can be achieved with a training set of logical queries and their corresponding answers, together with a partition of of the set of answers into disjoint preference sets $P^{+}$ and $P^{-}$. The training objective is then two-fold: (1) entities in $P^{+}$ should be ranked higher than those in $P^{-}$, and (2) all entities in $P^{+}\cup P^{-}$ (which satisfy the original logical query) should still rank above entities that do not meet the logical constraints of the query.
To achieve this, we adopt the RankNet loss (burges_learning_2005), which encourages pairs of relevant and non-relevant items to be ordered correctly. Given two entities $e_{i}$ and $e_{j}$ with scores $a[e_{i}]$ and $a[e_{j}]$, RankNet minimizes the binary cross-entropy over the probability $\sigma(a[e_{i}]-a[e_{j}])$ that $e_{i}$ should be ranked above $e_{j}$.
For NQR, we define two complementary loss terms:
$\mathcal{L}_{\text{NQR}}=\underbrace{\mathbb{E}_{(e^{+},e^{-})\sim(P^{+},P^{-})}\!\left[-\log\sigma\!\left(a[e^{+}]-a[e^{-}]\right)\right]}_{\text{(1) Preference-based ranking}}\;+\;\underbrace{\mathbb{E}_{(e^{+},e_{r})\sim(P^{+}\cup P^{-},\mathcal{N}_{m})}\!\left[-\log\sigma\!\left(a[e^{+}]-a[e_{r}]\right)\right]}_{\text{(2) Query-consistency ranking}},$ (12)
where $\mathcal{N}_{m}$ is a set of $m$ randomly sampled negative entities from $\mathcal{V}$ that are not valid answers to the query. The first term enforces the ordering implied by the preference set, while the second term ensures that preferred and non-preferred but valid answers remain ranked higher than irrelevant entities. We use gradient descent on this loss to update the parameters of the MLP in NQR.
6 Experiments
We aim to evaluate to what extent we can incorporate soft constraints as expressed in preference sets, while preserving overall query answering performance. We extend current benchmarks for query answering with preference data, and we run experiments comparing the performance of the Cosine and NQR updates, together with a baseline based on gradient-boosting approaches for learning to rank.
6.1 Datasets
Prior work on query answering on KGs has relied on datasets containing query-answer pairs $(q,\mathcal{A})$, where $q$ is a query and $\mathcal{A}\subset\mathcal{V}$ is the set of answers computed over a KG (ren_query2box_2020; ren_beta_2020). We extend these datasets by deriving preference data from observable clusters of entities in the set $\mathcal{A}$. This process consists of three steps (we provide additional details in Appendix  B):
1. Embedding textual descriptions. In order to capture soft similarity relationships that go beyond the structure of the graph and are harder to encode via logical constraints, we retrieve textual descriptions of entities in the KG and compute embeddings for each of them using a text embedding model.
Figure 2: Example dendrogram used to generate preference sets. Highlighted in green are preferred entities, other entities are non-preferred.
2. Clustering. We apply Hierarchical Agglomerative Clustering using cosine similarity as the distance metric. As shown in Fig. 2, the result is a dendrogram that iteratively merges the most similar entities into clusters, capturing similarities at different levels.
3. Answer set partitioning. We traverse the dendrogram from the root node towards the leaves, and if a cluster contains at least 20% of the entities in the answer set [^1] We define this limit to avoid clusters with too few entities., we use it to form a preference set. For each entity $e_{i}\in\mathcal{A}$, we define its preference label as $l_{i}=1$ if it is in the cluster, and $l_{i}=0$ otherwise. An example of a preference set is shown in Fig. 2, with entities labeled as $l_{i}=1$ are highlighted in green, and the rest of the entities are labeled as $l_{i}=0$. This partition forms the preference set $P(T)=\{(e_{1},l_{2}),\ldots,(e_{T},l_{T})\}$. We then denote an instance in our extended datasets as $(q,\mathcal{A},P(T))$, with instances having values of $T$ between 10 and 100.
Prior work has relied on subsets of encyclopedic KGs such as NELL (carlson_toward_2010) and Freebase (bollacker_freebase_2008) to benchmark methods for query answering. These KGs contain statements about people, organizations, and locations around the world. For our experiments, we select FB15k237 (toutanova_observed_2015) and the complex queries generated for it by ren_beta_2020. To further explore the potential of query answering in a different domain we use Hetionet (himmelstein_systematic_2017), a biomedical KG containing $\sim$2M edges between genes, diseases, and drugs, among other types. For Hetionet, we extract complex queries of different types, and then for both datasets we run the three steps above to collect preference data. We consider 14 types of first order logic queries of varying complexity, including negation. We present statistics of the final datasets and splits in Table 1. For training NQR, we only use the single type of 1-hop, link prediction queries. For the validation and test sets we use the full set of 14 types of complex queries. More detailed statistics can be found in Appendix  C.
Table 1: Statistics of the datasets used in our experiments.
Knowledge Graph Queries Preference Sets
Dataset Entities Relations Edges Train Validation Test Train Validation Test
FB15k237 14,505 237 310,079 8,472 24,435 24,505 42,360 122,175 122,525
Hetionet 45,158 24 2,250,198 55,772 24,427 24,494 278,860 122,135 122,470
6.2 Evaluation
We evaluate performance on the task of interactive query answering with soft constraints as follows: given an instance $(q,\mathcal{A},P(T))$ in the test set and a list of initial scores $a(q)$ computed by a base QA model, we select subsets of ${P}(t)\subseteq P(T)$ incrementally, with $1\leq t\leq 10$. For each subset we compute the adjusted score (Eq. 8), and then we evaluate the following:
1. Pairwise accuracy: Measures the extent to which preferred entities are ranked higher than non-preferred entities. Let $\texttt{rank}(e,a(q))$ be the position of entity $e$ when sorting the list of scores $a(q)$ in descending order. We define the pairwise accuracy as follows:
$\operatorname{PA}=\sum_{e^{+}\in P^{+}}\sum_{e^{-}\in P^{-}}\mathbbm{1}\left[\texttt{rank}(e^{+},a(q))<\texttt{rank}(e^{-},a(q))\right],$ (13)
where $\mathbbm{1}[\cdot]$ is an indicator function. Importantly, while score updates are computed using a subset of $P(T)$ of maximum 10 preferences, PA is computed using the full set, which can contain up to 100 preferences. This allows us to evaluate whether methods can generalize from a limited number of examples.
2. Ranking metrics: Following prior work on CQA, we evaluate evaluate query answering performance using ranking metrics. For an answer $e\in\mathcal{A}$ to a query, the reciprocal rank is $1/\text{r}(e,a^{(t)})$, and the Mean Reciprocal Rank (MRR) is the average over all queries [^2] We compute filtered ranking metrics, where other known answers to a query are removed before computing $\text{r}(e,a^{(t)})$.. Hits at $k$ (H@k) is defined by verifying if, for each answer $e$, its rank $\text{r}(e,a^{(t)})$ is less than or equal to $k$ (in which case H@k = 1, and H@k = 0 otherwise), and averaging over all queries. We compute these ranking metrics using the original and adjusted scores to determine any effect on overall query answering performance.
3. NDCG: We further compute the Normalized Discounted Cumulative Gain at $k$ (NDCG@k), which jointly captures the two criteria measured by PA and traditional ranking metrics: the relative ordering of preferred versus non-preferred answers, and their positions within the ranked list. We assign relevance scores of $0,1,2$ to non-answers, answers in $P^{-}$, and in $P^{+}$, respectively. The discounted cumulative gain and its normalized version are
$\mathrm{DCG@}k=\sum_{e\in A_{i}(q)}\frac{2^{\mathrm{rel}(e)}-1}{\log_{2}(j+1)},\qquad\mathrm{NDCG@}k=\frac{\mathrm{DCG@}k}{\mathrm{IDCG@}k},$ (14)
where $\mathrm{IDCG@}k$ is the DCG of the ideal ranking. An NDCG@k of 1 indicates that all preferred answers appear above non-preferred and non-answers.
6.3 Baselines
As an alternative to our proposed linear score update (Equation˜8), we consider a traditional Learning-to-Rank (LTR) baseline based on LightGBM (ke2017lightgbm), a gradient boosting framework implementing the LambdaRank objective (burges_learning_2005; burges_learning_2006). For each training query, the model receives as input three scalar features for every entity $e$: the original score $a[e]$, and the average positive and negative similarities, $\delta(P^{+},e)$ and $\delta(P^{-},e)$. Graded relevance labels are assigned following standard LTR practice (burges2010ranknet): entities in $P^{+}$ receive a label of 2, entities in $P^{-}$ receive 1, and a random set of 100 negative entities is assigned 0. The model is trained to predict ranking scores that order entities according to their relative relevance by directly optimizing NDCG via the LambdaRank objective. This baseline allows us to compare our proposed linear score update based on shifts from the original score, with a nonlinear approach that predicts the score directly using a nonlinear function of relevant features.
6.4 Experimental details
For all models, we find the best values of hyperparameters using the validation set via grid search. In all cases, we use QTO (bai2023qto) as the base QA model. More details about training and hyperparameter tuning can be found in Appendix D. We implement all of our experiments using PyTorch (paszke_pytorch_2019), and run them on a machine with one NVIDIA A100 GPU and 120GB of RAM.
7 Results
Figure 3: Results on interactive query answering with soft preferences, on FB15k237 and Hetionet.
Soft constraints and ranking quality.
We present results of pairwise accuracy, MRR, and H@10, and NDCG@10 in Fig. 3. For reference, we denote the metrics of the base QA model as Unconstrained, which allows us to determine how the adjustments in scores introduced by the methods we consider affect the performance of the original scores. We include 95% confidence intervals as shaded regions around the curves; these are narrow and often indistinguishable, indicating that the observed differences are consistent across queries.
Pairwise accuracy.
Across both datasets, we observe that all methods manage to achieve a pairwise accuracy significantly higher than the starting value (when no preferences are taken into account), which increases consistently as larger preference sets are provided. This indicates that all methods correctly adjust the scores in a way that entities in $P^{+}$ are ranked higher than those in $P^{-}$. Notably, the Cosine method achieves the highest values, which reaches a 87.7% in FB15k237, and 75.7% in Hetionet.
Ranking metrics.
The ranking metrics (MRR and H@1) behave differently in comparison with PA. We observe that when preference sets are small, with sizes between 1 and 4, score adjustments often cause a degradation in ranking quality with respect to the original scores. The fact that PA is already high at such sizes of preference sets indicates that the adjusted scores greedily place entities in $P^{+}$ above $P^{-}$, possibly by placing wrong answers between them, which harms ranking quality. Considering all methods, and especially in Hetionet, NQR causes the smallest drops in performance at small preference sets. As preference sets grow larger, the performance of all methods (with the exception of LightGBM, which struggles on FB15k237 and in H@10 on Hetionet) increases above the original. Both Cosine and NQR yield the highest ranking metrics, indicating that linear score updates are better for preserving ranking quality in comparison with the LightGBM baseline that predicts the new score directly.
NDCG.
The NDCG@10 results provide a unified view of the two underlying criteria: pairwise ordering and overall ranking quality. For small preference sets, all constrained models achieve higher NDCG@10 than the Unconstrained ranking, primarily due to better pairwise accuracy. Since both PA and ranking metrics improve as the size of the preference sets increases, NDCG@10 exhibits the same upward trend. Importantly, NDCG@10 allows comparing which method best satisfies the dual objective of query answering under soft constraints. We observe that for small preference sets, NQR performs best, suggesting that its learned conditional weights help preserve ranking quality while applying fine-grained score adjustments. For larger sets, the Cosine update surpasses NQR, matching or slightly improving ranking quality while achieving higher PA. This indicates that Cosine is a suitable general-purpose update rule, whereas NQR is preferable when fewer preferences are available.
We present additional results in Appendix˜E where metrics are categorized according to each of the types of complex queries in the datasets. The results show that in average Cosine performs best at balancing similarity constraints with global ranking performance, followed by NQR, which for certain query types, preserves best global ranking performance.
Figure 4: Trajectories of pairwise accuracy vs MRR for the Cosine update. Each subplot corresponds to a different value of $\alpha$, with lines colored by values of $\beta$. Highlighted circles indicate the initial interaction step (t=1), and the following points show increasingly larger preference sets.
Trade-offs induced by the Cosine update.
Figure˜4 illustrates how pairwise accuracy (PA) and mean reciprocal rank (MRR) evolve jointly for different values of $\alpha$ and $\beta$ in the Cosine update (Equation˜10), as the number of interaction steps increases. Across both datasets, increasing $\alpha$ (which implies placing more weight on the original scores), prevents significant drops in MRR, but also limits the maximum PA that can be reached. Smaller $\alpha$ values, on the other hand, allow stronger preference propagation, and the effect on MRR varies depending on the value of $\beta$. This behavior reflects the trade-off in the update: higher $\alpha$ yields conservative adjustments that favor stability, while lower $\alpha$ enables more aggressive score shifts guided by similarity signals.
Actionable control via $\beta$.
Within each $\alpha$ setting, varying $\beta$ determines the relative emphasis placed on positive versus negative preferences. Larger $\beta$ values (yellow curves) consistently increase PA with limited loss in MRR, while negative $\beta$ values (dark red) degrade both, indicating that over-penalizing negative examples distorts the global ranking. When $\alpha=0.25$ and after 10 interactions, the values closer to the Pareto front are $\beta=0.25$ and $\beta=0$ and $\beta=0.5$. Values of $\beta=0.5$ increase PA further, at the cost of lower MRR, and viceversa for $\beta=0$. This shows that $\beta$ offers an interpretable control mechanism: users prioritizing ranking fidelity can select lower $\beta$ values to protect MRR, whereas those seeking stronger preference alignment can increase $\beta$ to boost PA. This tunability makes the Cosine update practically useful for downstream applications, enabling domain experts to adjust the balance between preserving original rankings and enforcing similarity constraints according to their needs.
Qualitative example.
Table˜3 shows a 2-hop query from FB15k-237, where the task is to rank awards associated with people born in New York City. Applying similarity constraints on the target variable using the Cosine update sharply improves ranking quality: literary awards in the positive set are promoted into the Top-5, while unrelated film and TV awards drop in rank. NDCG@10 increases from 0.0 to 0.60, illustrating how a small preference set can adjust the ranking towards the soft constraint that awards returned as an answer to the query should be literary rather than film or music awards.
Table 2: Average execution time per query for different reranking methods on FB15k237.
Method Runtime (ms) $\Delta$ (ms)
Unconstrained 0.6
LightGBM 14.4 +13.8
Cosine 2.0 +1.4
NQR 1.8 +1.2
Runtime.
We present the average runtime in milliseconds per query in the FB215k237 dataset for all methods in Table˜2, where we include the added runtime ($\Delta$) with respect to the Unconstrained baseline. We note that LightGBM results in very high overhead, by adding 13.8 ms per query, due to the use of sequential tree traversal. On the other hand, the Cosine and NQR methods add very little overhead of up to 1.4 ms per query. While the average runtime of Cosine seems higher than NQR, a paired t-test does not show that the difference is statistically significant.
[…]
$q(v_{1},v_{2})=\texttt{PlaceOfBirth}(\texttt{New York City},v_{1})\wedge\texttt{NominatedFor}(v_{1},v_{2})$
$^{+} PP^{-}$
Hugo Award for Best Novel Grammy Award for Best Rock Instrumental Performance
World Fantasy Award for Best Novel Grammy Award for Best Country Song
World Fantasy Award for Best Novella Grammy Award for Best Country Performance
Initial Top-5 Top-5 after Cosine update
Golden Globe Award for Best Original Score $\blacktriangle$ Nebula Award for Best Novel
Academy Award for Best Original Song $\blacktriangle$ Hugo Award for Best Novella
Academy Award for Best Director $\blacktriangle$ Hugo Award for Best Short Story
Academy Award for Best Original Screenplay $\blacktriangle$ Hugo Award for Best Professional Artist
Nobel Prize in Literature $\blacktriangle$ Hugo Award for Best Novel
NDCG@10: 0.0 NDCG@10: 59.9
8 Conclusion
We have introduced the problem of interactive knowledge graph query answering with soft entity constraints. We formulate it as an extension that addresses the limitations of logical queries over knowledge graphs, which are restricted by the expressivity of first order logic and the schema of the graph.
We introduce efficient methods for incorporating soft constraints in complex queries. Since they reuse embeddings of a base CQA model, the methods are lightweight, requiring tuning only two parameters or a small MLP. In our experiments, we find that these methods are able to adjust query scores to capture preferences about their answers, while increasing ranking performance as more exemplary preferences are provided. In practice, we observe that our methods are fast, adding only up to 1.4 ms overhead to the base QA model.
With soft constraints, we explore a new and flexible way to interact with graph databases, which can allow users to specify their preferences by providing examples interactively, leading to query answering systems that can quickly adapt to user feedback.
Appendix A Appendix
Appendix B Generating preference data
In our experiments, we obtain preference data by clustering answers to queries over a KG using embedding models that map the description of an entity to an embedding of fixed dimension. In this section we provide additional technical details of this process.
Embedding textual descriptions.
We rely on text embedding models publicly available on Hugging Face [^3] https://huggingface.co/ for computing embeddings of entities based on their textual descriptions. For entities in FB15k237, we use NovaSearch/stella_en_400M_v5444https://huggingface.co/NovaSearch/stella_en_400M_v5, with an embedding dimension of 1,024. For Hetionet we employ Alibaba-NLP/gte-Qwen2-7B-instruct [^5] https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct, which at the time of writing is ranking 1st in the Massive Text Embedding Benchmark [^6] https://huggingface.co/spaces/mteb/leaderboard in the medical domain. The embedding dimension of this model is 3,584.
Clustering.
We execute hierarchical clustering using the average linkage method and cosine distance as the dissimilarity metric, applied to the target embeddings. This is implemented using the linkage function from the SciPy package. Average linkage is selected because it considers the mean pairwise distance between points in different clusters, which tends to produce more balanced and interpretable cluster structures. Cosine distance is used as it effectively captures angular similarity, which is more meaningful than Euclidean distance in high-dimensional embedding spaces, where the direction of vectors often encodes more relevant information than their magnitude.
Appendix C Dataset statistics
Figure 5: Types of complex queries used in our experiments.
We consider 14 types of complex queries over KGs introduced in prior work (ren_beta_2020), which include conjunctions, disjunctions, and negation. These queries can be represented using query graphs, which we illustrate in Fig. 5.
In our experiments we consider complex queries over the FB15k237 and Hetionet KGs, selecting those containing at least 10 answers and at most 100. For each query, we execute the algorithm for generating preference data (Appendix B) and generate 5 preference sets per query. The final number of queries and preference sets for each KG is shown in Table 4.
Table 4: Statistics of the datasets for query answering with soft preferences used in our experiments.
Split Structure 1p 2p 3p 2i 3i ip pi 2in 3in inp pin pni 2u up Total
FB15k-237
Train Queries 8,472 8,472
Preferences 42,360 42,360
Valid Queries 281 2,120 2,660 779 410 2,054 1,211 1,793 1,439 2,230 2,523 2,074 2,095 2,766 24,435
Preferences 1,405 10,600 13,300 3,895 2,050 10,270 6,055 8,965 7,195 11,150 12,615 10,370 10,475 13,830 122,175
Test Queries 327 2,101 2,717 843 500 1,942 1,184 1,805 1,569 2,257 2,460 2,031 2,052 2,717 24,505
Preferences 1,635 10,505 13,585 4,215 2,500 9,710 5,920 9,025 7,845 11,285 12,300 10,155 10,260 13,585 122,525
Hetionet
Train Queries 55,772 55,772
Preferences 278,860 278,860
Valid Queries 1,469 2,319 2,573 802 491 2,116 1,365 1,564 956 2,475 2,557 1,142 1,936 2,662 24,427
Preferences 7,345 11,595 12,865 4,010 2,455 10,580 6,825 7,820 4,780 12,375 12,785 5,710 9,680 13,310 122,135
Test Queries 1,472 2,309 2,562 850 556 2,087 1,376 1,556 1,004 2,473 2,535 1,137 1,938 2,639 24,494
Preferences 7,360 11,545 12,810 4,250 2,780 10,435 6,880 7,780 5,020 12,365 12,675 5,685 9,690 13,195 122,470
Appendix D Experimental details
Base query answering model.
We make use of QTO as the base query answering model bai2023qto. QTO relies on a neural link predictor, for which we use ComplEx (trouillon2016complex; lacroix2018complex_n3) trained with an auxiliary relation prediction objective (chen_relation_2021). We use an embedding dimension of 1,000, learning rate of 0.1, batch size 1,000, and a regularization weight of 0.05.
Cosine similarity.
We search for the best values of $\alpha_{p},\alpha_{n}$ via a grid search where each can take values in $\{0.1,0.25,0.5,0.75,0.9\}$, based on performance on the validation set. We define performance as the sum of pairwise accuracy and ranking performance.
NQR.
We describe the architecture of NQR in Table 5. When using the RankNet loss we select the best learning rate in $\{1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}\}$. For NQR, we perform a grid search where the learning rate takes values in $\{1\times 10^{-5},1\times 10^{-4}\}$, the margin in the preference loss in $\{0.05,0.1,0.25\}$, and the KL divergence weight in $\{0.1,1.0,10\}$.
Table 5: Layer-wise description of the model architecture in NQR. $t$ indicates the number of input preferences, which is variable. The embedding dimension is 1,000, and in specific layers we concatenate the score of an entity, indicated as +1+1.
Input Dimension Output Dimension
Preference embedding
Self-attention $(t,1000+1)$ $(t,1001)$
LayerNorm $(t,1001)$ $(t,1001)$
Linear + ReLU (fc1) $(t,1001)$ $(t,1000)$
Mean-pooling $(t,1000)$ 10001000
Score adjustment
Linear + ReLU $(2000+1)$ 10001000
Linear + tanh 10001000 11
Appendix E Additional results
In our main experiments, we report the performance of different methods when presented with incremental feedback with preference sets of size 1 up to 10, averaged over all queries. This requires queries that have at preference set (combining both preferred and non-preferred entities) of at least 10 entities, such that we can evaluate methods over the entire range of preference set sizes.
We present here results over all types of complex queries we consider in our experiments. As in our main experiments, we compute the metrics for preference sets of increasing sizes, from 1 up to 10. Here, we report the metrics averaged over all 10 steps. We present results categorized by query structure (illustrated in Fig. 5) in Tables 6 and 7.
We observe that in average, Cosine and NQR result in the best results, with Cosine achieving a good balance between pairwise accuracy and global ranking performance, as shown by the results in NDCG@10. The results by query type also indicate cases where all models struggle to maintain the ranking performance of the Unconstrained model: in the FB15k237 dataset and for ip and pi queries, all models result in lower MRR, with NQR being the closest. Similarly, in Hetionet all models underperform the Unconstrained baseline in 2u queries, though NQR achieves a close value.
Table 6: Per query type results of interactive query answering with feedback on the FB15k237 dataset.
Method 1p 2p 3p 2i 3i ip pi 2in 3in inp pin pni 2u up avg
Average Pairwise Accuracy
Unconstrained 58.44 55.67 55.30 55.55 53.45 56.19 54.97 58.06 54.99 53.41 54.40 58.70 59.08 55.74 56.00
LightGBM 80.10 66.07 59.85 71.98 66.97 61.55 61.70 74.45 69.43 61.25 62.89 73.14 75.09 67.58 68.00
Cosine 85.66 85.06 85.80 81.42 77.23 84.68 82.70 82.99 81.64 84.61 84.45 83.60 83.54 83.99 83.38
NQR 76.80 76.11 77.45 71.75 68.76 76.56 74.09 73.78 71.99 74.66 75.29 74.34 73.71 74.27 74.26
Average MRR
Unconstrained 19.97 16.32 15.73 24.02 27.73 21.90 22.27 11.32 15.39 11.55 9.53 3.88 16.08 16.44 16.58
LightGBM 16.93 15.34 14.69 19.82 24.25 18.90 19.31 11.81 15.18 12.03 10.85 8.96 13.28 14.96 15.45
Cosine 24.09 16.65 14.52 23.25 27.78 16.77 17.51 15.08 16.31 14.47 13.02 7.89 17.88 16.12 17.24
NQR 25.37 19.08 16.28 26.22 29.17 19.15 20.34 14.60 17.79 15.98 13.11 5.63 18.37 18.84 18.57
Average NDCG@10
Unconstrained 39.92 36.79 36.94 51.40 54.67 44.44 48.28 33.56 43.10 29.86 28.28 10.86 43.37 39.88 38.67
LightGBM 45.62 43.73 42.06 53.42 56.09 48.13 51.49 43.07 49.81 39.32 38.45 36.44 45.42 46.72 45.70
Cosine 58.64 55.67 53.23 62.01 63.42 56.37 56.52 54.77 56.52 53.36 51.76 33.15 60.11 56.84 55.17
NQR 57.07 56.39 54.71 61.95 62.79 57.66 58.33 49.37 55.30 51.77 48.81 20.87 56.70 57.00 53.48
Table 7: Per query type results of interactive query answering with feedback on the Hetionet dataset.
Method 1p 2p 3p 2i 3i ip pi 2in 3in inp pin pni 2u up avg
Average Pairwise Accuracy
Unconstrained 51.75 50.73 50.55 51.78 51.19 51.00 51.06 48.03 41.51 49.99 48.84 51.73 52.62 50.58 50.10
LightGBM 79.18 65.73 63.76 75.71 75.20 64.63 69.48 71.17 68.62 63.37 64.24 74.14 74.24 67.16 69.76
Cosine 75.27 73.30 71.65 72.13 71.48 73.19 73.57 74.19 74.58 73.29 71.12 75.06 70.70 69.21 72.77
NQR 68.17 63.19 60.78 65.96 65.30 65.40 67.12 66.84 69.39 61.60 60.67 67.28 63.44 59.59 64.62
Average MRR
Unconstrained 25.91 4.83 7.86 21.74 26.20 7.98 15.00 13.27 8.53 4.25 3.80 3.27 18.89 5.69 11.95
LightGBM 13.37 8.47 11.31 11.42 14.28 10.26 11.47 9.26 7.87 8.19 7.23 9.87 9.19 7.32 9.97
Cosine 22.07 9.26 9.66 20.40 24.45 10.61 13.94 13.52 12.75 8.77 8.30 9.24 15.06 8.23 13.30
NQR 26.39 8.87 9.32 22.89 27.23 10.77 15.51 15.11 13.20 7.82 7.66 7.23 18.72 8.22 14.21
Average NDCG@10
Unconstrained 44.20 19.83 26.34 43.06 45.51 24.72 38.31 33.33 22.95 19.29 17.58 9.62 48.75 23.92 29.81
LightGBM 41.45 41.17 49.29 39.12 41.06 43.87 40.70 35.96 29.84 43.06 39.60 38.24 40.35 43.64 40.52
Cosine 53.28 45.09 49.16 53.79 56.23 46.80 47.95 46.59 44.64 46.79 44.57 35.94 52.21 47.01 47.86
NQR 53.04 42.28 41.44 53.16 55.79 45.77 48.91 45.97 43.34 40.97 40.23 27.65 54.24 43.83 45.47